OpenStudy (anonymous):

Select two equations that can represent the system 3y+2x=9 x-y=2 in terms of the y-variable.

OpenStudy (anonymous):

These are multiple choice answers and I have to pick 2

OpenStudy (anonymous):

A. 3y+2(y+2)=9 B. (-3y+9)-y=2 C. 2y=-3y+9 D.(-3/2y+9/2)-y=2

OpenStudy (anonymous):

I'll give a medal help fast

OpenStudy (texaschic101):

x - y = 2 x = y + 2 3y + 2x = 9 3y + 2(y + 2) = 9 <== this is one (-3/2y + 9/2) - y = 2 <==this is one

OpenStudy (anonymous):

Oh ok thank you